Suppose the short-run cost function of Tyrell Corporation is 𝒄 = πŸ“πŸŽπŸŽπŸŽ + πŸπŸŽπŸŽπ’’ + πŸ–π’’πŸ.Β 

a.Β Β Find the level/s of Β that satisfies/satisfy the first-order and second-order conditions for the minimisation

of Tyrell’s average cost (𝒂𝒄). [4]

Show that 𝒂𝒄 equals marginal cost (π’Žπ’„) at the level/s of Β obtained in (a). [2

1
Expert's answer
2021-06-07T17:02:19-0400

(a)

Given cost function:

"C = 5000 + 100q + 8q^2"

Average cost (AC)

"=\frac{ C}{q}"

"=\frac{5000}{q}+100+8q"

To minimize AC, FOC :-

"=\frac{\delta(AC)}{\delta q}=0"

or, "5000(-1)q^{-2} + 8 = 0"

or, "q = 25"


(b)

SOC :-

"\frac{d^{2}(AC)}{dq^{2}}=-5000(-2)q^{-3}"

Substituting q = 25 we get 0.64, which is > 0. Hence SOC is satisfied. AC must be minimum at q = 25.

Substituting q=25 in AC we get 500.

Now, MC:-

"\frac{\delta C}{\delta q}=100+16q"

Substituting q=25 in MC we get 500.

Hence MC = AC at q=25.